For a birthday party, Marissa bought three cartons of ice cream at $4.59 each and two packages of plastic bowls at $3.00 each. If food is not taxed and non-emergency grocery items are taxed at the rate of 5%, what was her total bill?
Answer by rfer(16322)

3*4.59+2*3*1.05=13.77+6.30=$20.07